# Can Language Models Verbalize Well-Defined Notions of Uncertainty?

### v2

@@ 4.1 Uncertainty Distillation on LLMs
We validate uncertainty distillation on these datasets using two modern instruction-tuned LLMs, Llama-3.2-3B-Instruct Dubey et al. (2024) and Ministral-8B-Instruct-2410 Jiang et al. (2023). When performing uncertainty distillation with Ministral-8B, we use LoRA Hu et al. (2021).
Baselines
For the Lexical baseline, we extract token-level probabilities from the language model on our training/calibration split [^7] As we do not have an initial fine-tuning step, these are equivalent. and use this to train an isotonic regression model to calibrate the average token-level probability for each answer. [^8] We use the average probability rather than the sequence probability to normalize over different lengths, as Kuhn et al. (2023) find this improves performance. Prompting: To measure the model’s ability to verbalize its confidence prior to uncertainty distillation we prompt the model to output its own confidence in its answer without any additional training. We report this baseline for these models, and discuss the prompts used in Appendix C. We also compare to P(IK) from Farquhar et al. (2024) which learns a mapping from hidden states to uncertainty scores, and P(True) from Kadavath et al. (2022). Finally, we compare to the Semantic Clustering approach from Farquhar et al. (2024). Both P(True) and Semantic Clustering generate 20 samples from the model to compute uncertainty scores, unlike our approach which uses a single generation.
4.2 Targeted analysis
In §missing 3, we assume that we have access to held-out calibration data. However, due to the opacity and size of pretraining datasets, it is increasingly challenging to guarantee that this assumption holds. We therefore test uncertainty distillation in a setting where we can know with certainty whether the calibration set is in a model’s pretraining data. We examine this question and the impact of adding varying numbers of incorrect answers during uncertainty distillation in §missing 6.
Dataset
In this setting, we use the Super-NaturalInstructions dataset (SNI; Wang et al., 2022). We select $15$ English Q&A tasks with short-form answers. We focus on Q&A tasks for which a single correct answer exists (e.g. multiple choice problems, short-form span extraction, math problems, etc.) and thus for which correctness of a model’s prediction can reliably and efficiently be computed after normalizing lexical forms without resorting to methods such as LLM verification. We use 1,000 samples to obtain our Monte Carlo estimate of confidence (see Appendix B for details on how number of samples affects successful confidence estimation).
Models
We perform uncertainty distillation on FLAN-T5 Chung et al. (2022), an instruction-tuned model trained on a dataset containing the SNI tasks. Importantly, we not only verify that Flan-T5 has been instruction-tuned on our tasks, but has seen samples from the calibration set of our test tasks. This allows us to investigate the effect of data contamination on calibration of verbalized confidences.
To construct a similar model which has not seen our calibration data, we instruction-tune a T5-Large model on a remaining subset of the English tasks in the SNI dataset, making sure to explicitly hold out the $15$ tasks we use in our uncertainty distillation experiments. The result is an instruction-tuned model which we refer to as Instruct-T5, capable of performing our target Q&A tasks without having seen these tasks during training. In other words, the samples we obtain from this model do not require Instruct-T5 to be pre-trained on that specific task. See Appendix E for more details on our data selection and instruction-tuning. We train and evaluate uncertainty distillation on the combined dataset of these tasks and report the performance over the metrics described in §missing 4.3.
@@ 4.2 Targeted analysis
4.3 Metrics
We report the area under the receiver operating characteristic curve (AUROC), [^9] Calculated using scikit-learn 1.5.2 which represents the probability that a randomly chosen correct answer will be in a higher-confidence bin than a randomly chosen incorrect answer. This metric is well established in previous literature (see e.g., Hu et al. (2023)), and compares the relative rather than absolute probabilities, which allows us to use it effectively with discrete verbalized confidences. [^10] We do not report Expected Calibration Error (ECE) (Nixon et al., 2019), as it requires comparing a continuous score to the prediction’s true label, while we have five discrete bins without an obvious corresponding numerical score. Furthermore, it requires the choice of several hyperparameters which can have a large impact on performance. Baseline methods that return a continuous score are binned to five categories to represent converting to a comparable verbalized confidence. For all methods, we plot the percentage of accurate answers in each bin to examine if confidence corresponds well with accuracy. We also report overall model accuracy, to evaluate the tradeoff between accuracy and calibration. Finally, we report high accuracy, accuracy of high-confidence predictions (in “very high” and “high” bins). As an established use-case for verbalized confidences is to reject lower-confidence predictions, this provides information about how useful the LLM’s predictions are. [^11] The fact that high accuracy is not perfect also highlights a risk of confidence estimation: namely, that it increases trust in an answer that still may be incorrect.
5 Results and discussion
Figure 2 shows some of our results comparing uncertainty distillation to the lexical uncertainty baseline in terms of average accuracies in each confidence bin [^12] We present the remaining two settings in Appendix F.. In plots like this, the goal is to observe a diagonal trend line where outputs reported to have high confidence indeed have high accuracy, and those in the low confidence bins have lower accuracy. We find that the verbalized confidences produced by uncertainty distillation are highly interpretable, with high correspondence between accuracy of answers within a bin and that bins confidence. In contrast, confidence scores generated by the baselines may not correspond well with the actual accuracies within that bin. For instance, accuracy within the lowest confidence bin for the prompting baseline is 0.684 with Llama-3B on SocialIQA, while accuracy within the highest confidence bin is 0.651.
Table 1 summarizes these plots in terms of AUROC score. AUROC is consistently higher with uncertainty distillation than with other methods. We conclude that uncertainty distillation is an effective method for estimating confidence in an answer. AUROC is consistently highest for uncertainty distillation with both models on both datasets. In particular, we note that uncertainty distillation consistently achieves higher AUROC than semantic clusteringKuhn et al. (2023), despite semantic clustering requiring 20 samples and a computationally intensive clustering step at inference time: for instance, uncertainty distillation achieves AUROC of 0.784 with Llama-3B on SocialIQA, while semantic clustering achieves AUROC of 0.601.
The table also reports the accuracy of the highest confidence bin and the overall accuracy across all bins. While AUROC is the main metric for assessing calibration performance, accuracy is also useful for understanding the nuances of the result. We find that uncertainty distillation does not lead to notable drops in overall accuracy, and that accuracy in the highest bins increases dramatically. Uncertainty distillation achieves the best high accuracy in all but one case, Ministral-8B on MMLU. In this case, semantic clustering achieves high accuracy of 0.821 compared to uncertainty distillation’s high accuracy of 0.766. However, we again note the comparative computational complexity of semantic clustering, as well as the fact that uncertainty distillation’s high accuracy is significantly higher than the next highest, prompting with 0.643.
6 Analysis
@@
6.1 Assumption of calibration set
We compare the performance of FLAN-T5, which has been instruction-tuned on the calibration set, with the performance of Instruct-T5, which has not, in Table 2. We find that while uncertainty distillation still produces meaningful confidence bins for FLAN-T5, it no longer outperforms lexical uncertainty. We conclude that uncertainty distillation works in the absence of held-out calibration data, but not as effectively as token-level probabilities, which are likely well-calibrated due to the model’s previous training on these examples. We discuss results for these two models further in §missing 6.2 and Appendix D, and find that the behavior of FLAN-T5 differs significantly from results on models where we have an unseen calibration set.
6.2 Adding incorrect examples
@@
7 Conclusion
Findings
We find that uncertainty distillation leads to improved estimates of uncertainty in comparison to many strong baselines, including baselines that require considerably more samples at inference-time. One high-level takeaway is that with small models there appears to be a tradeoff between an LLM’s ability to predict its own confidence and overall model accuracy, but that this effect is less obvious with increasing model sizes. In our small-scale analysis, interventions that improve AUROC decrease accuracy and vice versa; however, with larger models we do not note as noticeable a decrease in accuracy compared to our baselines. Overall, we view our contribution as a significant step towards LLMs that can reliably reason about uncertainty, without requiring any auxiliary models or incurring additional inference-time compute.
Future work
While we focus on QA tasks, our method could be applied to tasks outside simple QA through the use of LLM verifiers to calculate binary correctness. Future work may also investigate the robustness of the model’s internal representation of uncertainty to domain shifts—do the confidence scores remain interpretable and well-calibrated on similar unseen tasks without further training? Looking beyond these immediate questions, LLMs that are able to verbalize meaningful confidences, for example thanks to our method, may be useful in a variety of applications requiring reasoning about uncertainty, such as medical diagnosis.
Limitations
Our experiments focus on established QA tasks which admit straightforward ways to assess correctness. In principle, our approach generalizes to more complex tasks involving longer-form generations, for example using an LLM verifier to establish correctness; we leave it as future work to experiment in these settings. Separately, the proposed approach may be useful in cases where a single generation involves multiple distinct claims that each need to be associated with distinct confidences. Future work should identify appropriate datasets to evaluate multi-claim uncertainty estimation. Finally, our experiments do not include models larger than 8 billion parameters due to compute limitations; however, we hope that our findings will encourage further study into uncertainty distillation for larger LLMs and in more general settings.
Appendix A Uncertainty distillation on supervised fine-tuned models
@@
We perform these experiments using the SQuAD benchmark (Rajpurkar, 2016). This is a machine-reading task where each question consists of a passage of text and one or more associated questions, each of which is answerable based on the text itself. As the test set has not been publicly released, we use the splits proposed by Du et al. (2017), which divides the publicly available available training and validation splits into train, test, and validation splits. We consider the first 60,000 examples in the training set to be training data, and the remainder to be our calibration set.
@@
We apply uncertainty distillation to T5-base (Raffel et al., 2020) finetuned on a portion of SQUAD. We use defaults for most hyperparameters, and report hyperparameters in Appendix I.
@@ Results
Appendix B Number of samples
Our Monte Carlo estimation of probability requires sampling repeatedly from a model before normalizing and calculating probability. In Figure 3, we show that the number of samples used to estimate the initial probabilities has a significant impact if chosen to be too low; however, there are diminishing returns as the number of samples increases. We therefore choose to use 1,000 samples in all of our experiments with FLAN-T5 and Instruct-T5, as more than that is unlikely to achieve anything but marginal improvement.
@@ Appendix B Number of samples
Appendix C Prompts
C.1 Mistral, Llama
Prompt baselines, uncertainty distillation
Answer the following question and state confidence in the answer (very low, low, medium, high, very high). Enclose concise reasoning in <reasoning> </reasoning> tags, confidence in <confidence> </confidence> tags, and your FINAL answer in <answer> </answer> tags without any of your work, like this: "If each of Lisa’s 7 chickens lays 6 eggs, how many eggs does Lisa have?
@@ C.1 Mistral, Llama
Answer the following question. Enclose concise reasoning in <reasoning> </reasoning> tags and your FINAL answer in <answer> </answer> tags without any of your work, like this: "If each of Lisa’s 7 chickens lays 6 eggs, how many eggs does Lisa have?
@@
C.2 Instruct-T5, FLAN-T5
@@
Appendix D Effects of post-hoc calibration on well-calibrated models
If the model’s initial predictions are poorly calibrated, the post-hoc calibration step should help to better align probabilities in the training data with the true likelihood of success; indeed, we find that post-hoc calibration has a significant positive effect with our largest models. For instance, Llama-3B on SocialIQA achieves 0.784 AUROC when trained on post-hoc calibrated data, and only 0.680 when identically trained on data without post-hoc calibration.
However, how does post-hoc calibration impact the model when the model is already well-calibrated on the specific task, or when the model has previously been trained on the calibration data? Figure 4 shows the reliability diagrams for T5-base on SQUAD and Instruct-T5 on SNI. The models’ predicted confidences align well with their actual accuracies; this allows us to investigate whether post-hoc calibration has a significant impact on AUROC. Additionally, FLAN-T5 has been previously tuned on our calibration set; this gives us a setting to investigate the impact of post-hoc calibration when unseen calibration data is unavailable.
In Table 5, we show the results of the smaller models trained with and without this post-hoc calibration step. We find no apparent benefit of post-hoc calibration for Instruct-T5 or fine-tuned T5-base. These models are already well-calibrated on their domains; therefore, a post-hoc calibrator does not significantly alter the output probabilities.
In the case of FLAN-T5, post-hoc calibration decreases AUROC. This suggests that in cases when unseen calibration data cannot be obtained, uncertainty distillation may be more effective without the post-hoc calibration step.
@@ Appendix D Effects of post-hoc calibration on well-calibrated models
Appendix E SuperNatural-Instructions Tasks
E.1 Target Calibration Tasks
As we describe in §missing 3, in this work we rely on the assumption that our target-tasks have a correct answer, in the sense that it can be easily verified that an answer is right or wrong. Although this is not a strict necessity for calibration, it allows for us to define our buckets in terms of expected accuracy, rather than e.g. an expected score. We therefore focus on short-form Q&A tasks, question-answer pairs whose answers consist of either selection from a fixed answer set (e.g. multiple choice or fixed choice) or single-word answers. We identify $15$ tasks from the SuperNatural-Instructions dataset (Wang et al., 2022) that fit our criteria, and hold out these tasks as our uncertainty prediction tasks.
@@ E.1 Target Calibration Tasks
E.2 Instruction-Tuning Tasks
Because most modern instruction-tuned models are trained on all of Super-NaturalInstructions, they have seen the our calibration target tasks during instruction-tuning. Therefore, we instruction-tune our own T5 model to test the effectiveness of our method on unseen tasks. Our model is trained on a subset of the SuperNatural-Instructions dataset (Wang et al., 2022). Specifically, we instruction-tune on the English split used in the original paper but we take out our target calibration tasks identified in §missing E.1. This gives us a training dataset of $879$ instruction-tuning tasks, with a total of roughly 1.2M training samples total.
@@ E.2 Instruction-Tuning Tasks
Appendix F Ministral plots
In Figure 5 we display the plots with Ministral-8B. As reflected in the AUROC score in Table 1, calibration is slightly worse; however, compared to baselines, it still does a more accurate job of forecasting confidence.
@@ Appendix F Ministral plots
Appendix G Instruction-Tuning T5
@@
We follow a standard recipe for instruction-tuning T5-Large, established in Wang et al. (2022). Specifically, we tune the model for 3 epochs with a batch size of $16$ and a learning rate of $5\times 10^{-3}$. We use the AdamW optimizer, and a constant learning rate schedule after a warmup period of $500$ steps. During instruction-tuning, we train the model with the semantic definition of each task prepended to the task input, and we similarly prompt the model when performing our target Q&A tasks.
Appendix H Resource Reporting
H.1 Compute Resources
@@
Here we report the compute resources used in this work. Instruction-tuning T5 took a total of $200$ GPU hours across $4$ NVIDIA-V100s. Running uncertainty distillation on Instruct-T5 and FLAN-T5 took 16 hours per model on a single NVIDIA-H100. Finetuning T5-base on SQUAD for our initial model took 3 hours on a single NVIDIA RTX 2080, and training using uncertainty distillation took 8 hours on a single NVIDIA-V100. Finetuning Ministral-8B (LoRA) and finetuning Llama-3B each took took three hours on two NVIDIA-A100s. Our lexical baseline for SQUAD took one hour on one NVIDIA RTX 2080; for SNI took three hours on one NVIDIA RTX 2080; for MMLU took three hours on one NVIDIA-A100; and for SocialIQA took thre hours on one NVIDIA-A100. Prompting for MMLU and prompting for SocialIQA took 1 hour on one NVIDIA-A100. Sampling for SQUAD took a total of 60 GPU hours on NVIDIA-V100s; for SNI took 45 GPU hours on NVIDIA-A100s; for SocialIQA took 350 hours on NVIDIA-A100s; and for MMLU took 350 hours on NVIDIA-A100s.
H.2 Resource Intended Use
@@
Super-NaturalInstructions (SNI) is an open-source instruction tuning dataset, released under the Apache License. [^13] Available here: https://github.com/allenai/natural-instructions The intended use of SNI is to instruction-tune language models to learn to follow instructions, and to evaluate a model’s ability to follow instructions on unseen tasks. While we use the SNI dataset for precisely this purpose during instruction-tuning, we also use $15$ held-out tasks to serve as uncertainty quantification tasks. This does not necessarily fall under the intended use of instruction-tuning; however, the authors of SNI also mention that the dataset may serve as a large, multi-task natural language resource (Wang et al., 2022), and our usage of the target calibration tasks does fall under this use case.
@@
The Stanford Question Answering Dataset (SQUAD) Rajpurkar (2016) is distributed under the Creative Commons Attribution-Sharealike 4.0 license, which permits use of the dataset as long as it is properly attributed and as long as the results are distributed under the same license. As we cite the paper and plan to publically release our code and models after acceptance, our use of this dataset is permitted under this license.
@@
SocialIQA Sap et al. (2019) is not explicitly licensed, but they state that they “ establish Social IQa as a resource” for future models.
@@
MMLU Hendrycks et al. (2020a) is published under the MIT license, which allows users to freely copy, use, and change the licensed material.
Appendix I Uncertainty distillation hyperparameters
@@
Model Epochs Learning rate Batch size Grad accumulation steps
@@
Table 7: Hyperparameters for training all T5 models but Instruct-T5 (see Appendix G for details). All models are trained with the AdamW optimizer.
@@
     $b\leftarrow$ bin$(p)$
     $z\leftarrow$ verbalize_confidence_map$(\hat{y},b)$

### v3

@@ B.1 Experimental setting
We perform uncertainty distillation on FLAN-T5 (chung2022scalinginstructionfinetunedlanguagemodels), an instruction-tuned model trained on a dataset containing the SNI tasks. Importantly, we not only verify that Flan-T5 has been instruction-tuned on our tasks, but has seen samples from the calibration set of our test tasks. This allows us to investigate the effect of data contamination on calibration of verbalized confidences.
To construct a similar model which has not seen our calibration data, we instruction-tune a T5-Large model on a remaining subset of the English tasks in the SNI dataset, making sure to explicitly hold out the $15$ tasks we use in our uncertainty distillation experiments. The result is an instruction-tuned model which we refer to as Instruct-T5, capable of performing our target Q&A tasks without having seen these tasks during training. In other words, the samples we obtain from this model do not require Instruct-T5 to be pre-trained on that specific task. See Appendix H for more details on our data selection and instruction-tuning. We train and evaluate uncertainty distillation on the combined dataset of these tasks and report the performance over the metrics described in § 4.3.
@@
B.2 Results
Assumption of calibration set
We compare the performance of FLAN-T5, which has been instruction-tuned on the calibration set, with the performance of Instruct-T5, which has not, in Table 5. We find that while uncertainty distillation still produces meaningful confidence bins for FLAN-T5, it no longer outperforms lexical uncertainty. We conclude that uncertainty distillation works in the absence of held-out calibration data, but not as effectively as token-level probabilities, which are likely well-calibrated due to the model’s previous training on these examples. We discuss results for these two models further in § B.2 and Appendix G, and find that the behavior of FLAN-T5 differs significantly from results on models where we have an unseen calibration set.
@@ B.2 Results
Adding incorrect examples
@@
B.3 Analysis
One high-level takeaway is that with small models there appears to be a tradeoff between an LLM’s ability to predict its own confidence and overall model accuracy, but that this effect is less obvious with increasing model sizes. In our small-scale analysis, interventions that improve AUROC decrease accuracy and vice versa; however, with larger models we do not note as noticeable a decrease in accuracy compared to our baselines. Functionally, UD combines aspects of two tasks: the model’s original question answering ability and uncertainty quantification. Large models are both less prone to catastrophic forgetting(ramasesh2022effect)and more effective at multitask learning than smaller models(chung2022scalinginstructionfinetunedlanguagemodels). With this framing, the fact that larger models’ accuracy is less impaired by the finetuning process of uncertainty distillation indicates that model scale plays a significant role in an accuracy/performance tradeoff, and increasing model scale or training in an explicitly multi-task setting may decrease the likelihood of drops in accuracy.
Appendix C Uncertainty distillation on supervised fine-tuned models
@@
We perform these experiments using the SQuAD benchmark (rajpurkar2016squad). This is a machine-reading task where each question consists of a passage of text and one or more associated questions, each of which is answerable based on the text itself. As the test set has not been publicly released, we use the splits proposed by DBLP:journals/corr/DuSC17, which divides the publicly available available training and validation splits into train, test, and validation splits. We consider the first 60,000 examples in the training set to be training data, and the remainder to be our calibration set.
@@
We apply uncertainty distillation to T5-base (raffel2020exploring) finetuned on a portion of SQUAD. We use defaults for most hyperparameters, and report hyperparameters in Appendix L.
@@ Results
Appendix D Number of samples
Our Monte Carlo estimation of probability requires sampling repeatedly from a model before normalizing and calculating probability. In Figure 3, we show that the number of samples used to estimate the initial probabilities has a significant impact if chosen to be too low; however, there are diminishing returns as the number of samples increases. We therefore choose to use 1,000 samples in all of our experiments with FLAN-T5 and Instruct-T5, as more than that is unlikely to achieve anything but marginal improvement. For the larger models, we select 100 samples, as this appears to be the elbow of the curve in Figure 3, and as sampling 1000 samples from the large models would be computationally prohibitive. We note, however, that based on these results, this hyperparameter can be changed to improve efficiency or effectiveness of the method as is required by each task.
@@ Appendix D Number of samples
Appendix E Prompts
E.1 Mistral, Llama
Prompt baselines, uncertainty distillation (MC)
Answer the following question and state confidence in the answer (very low, low, medium, high, very high). Enclose concise reasoning in <reasoning> </reasoning> tags, confidence in <confidence> </confidence> tags, and the letter of your FINAL answer in <answer> </answer> tags without any of your work, like this: "If each of Lisa’s 7 chickens lays 6 eggs, how many eggs does Lisa have?
@@ E.1 Mistral, Llama
Prompt baseline, uncertainty distillation (open)
"You are a helpful AI assistant. Answer the following math question as briefly as possible and accurately. Enclose confidence in the answer (very low, low, medium, high, very high) after the answer in <confidence> </confidence> tags, like so: <confidence> very high </confidence>."
@@
Answer the following question. Enclose concise reasoning in <reasoning> </reasoning> tags and the letter of your FINAL answer in <answer> </answer> tags without any of your work, like this: "If each of Lisa’s 7 chickens lays 6 eggs, how many eggs does Lisa have?
@@
Sampling, lexical baseline (open)
"You are a helpful AI assistant. Answer the following math question as briefly as possible and accurately."
E.2 Instruct-T5, FLAN-T5
@@
E.3 LLM-as-a-judge
We are evaluating answers to the question {̈question}\̈\ Here are two possible answers:
Possible Answer 1: {text1}
Possible Answer 2: {text2}
Is Possible Answer 1 equivalent to Possible Answer 2, or do the answers contradict? Respond only with ’equivalent’ or ’contradictory’.
Appendix F Bin and label choice
In the main experiments, we examine the effect of UD with five bins and a verbalized naming scheme. However, in Figure 4, we examine the effect of running UD on SocialIQA with Llama-3B while varying the number of bins (and thus necessarily changing the labeling scheme). Here, we find appropriate calibration regardless of number of bins. Notably, even changing the labeling scheme to numerical percentages does not result in a change in performance, suggesting that UD is robust to variance in labeling schemes. We use five bins as the default, as it offers enough bins to be challenging while avoiding the problems of sparsity (and thus noisiness) in bins that arise with larger bin sizes.
Figure 4: Running UD on SIQA with Llama-3B and changing the number of the bins, or the labeling scheme, has no noticeable effect on the efficacy of UD aside from increased sparsity in bins. As in other figures, bins with fewer than 10 samples are not plotted.
Appendix G Effects of post-hoc calibration
If the model’s initial predictions are poorly calibrated, the post-hoc calibration step should help to better align probabilities in the training data with the true likelihood of success. In Table 8, we compare the miscalibration of the training data (measured through ECE with 30 bins) to the performance of models with and without post-hoc calibration during data generation. Unsurprisingly, we find that post-hoc calibration has positive effect corresponding to the initial miscalibration of the training data. For instance, Llama-3B on SocialIQA achieves 0.784 AUROC when trained on post-hoc calibrated data, and only 0.691 when identically trained on data without post-hoc calibration, with an ECE of 0.10. However, Ministral-8B on SocialIQA has a comparatively small ECE of 0.026, and the performance without post-hoc calibration is equivalent to the performance with post-hoc calibration. We conclude that the decision to include post-hoc calibration can be quickly and cheaply made by simply measuring the calibration of the annotated training data.
Dataset Model With Post-hoc No Post-hoc Training Data ECE
MMLU Ministral-8B 0.693 0.689 0.033
Llama-3B 0.743 0.714 0.039
SIQA Ministral-8B 0.671 0.673 0.026
Llama-3B 0.784 0.691 0.100
Table 8: AUROC of large models with and without post-hoc calibration at training time. We find that post-hoc calibration tends to improve performance, most dramatically with Llama-3B on SIQA.
We further analyze how post-hoc calibration impacts the model when small models are already well-calibrated on the specific task. Figure 5 shows the reliability diagrams for T5-base on SQUAD and Instruct-T5 on SNI. The models’ predicted confidences align well with their actual accuracies; this allows us to investigate whether post-hoc calibration has a significant impact on AUROC for smaller models. Additionally, FLAN-T5 has been previously tuned on our calibration set; this gives us a setting to investigate the impact of post-hoc calibration when unseen calibration data is unavailable and the model is presumably correctly confident in its predictions.
In Table 9, we show the results of the smaller models trained with and without this post-hoc calibration step. We find no apparent benefit of post-hoc calibration for Instruct-T5 or fine-tuned T5-base. These models are already well-calibrated on their domains; similarly to large models, a post-hoc calibrator does not significantly alter the output probabilities.
In the case of FLAN-T5, post-hoc calibration decreases AUROC. This suggests that in cases when unseen calibration data cannot be obtained for small models, uncertainty distillation may be more effective without the post-hoc calibration step.
@@ Appendix G Effects of post-hoc calibration
Appendix H SuperNatural-Instructions Tasks
H.1 Target Calibration Tasks
As we describe in § 3, in this work we rely on the assumption that our target-tasks have a correct answer, in the sense that it can be easily verified that an answer is right or wrong. Although this is not a strict necessity for calibration, it allows for us to define our buckets in terms of expected accuracy, rather than e.g. an expected score. We therefore focus on short-form Q&A tasks, question-answer pairs whose answers consist of either selection from a fixed answer set (e.g. multiple choice or fixed choice) or single-word answers. We identify $15$ tasks from the SuperNatural-Instructions dataset (wang-etal-2022-super) that fit our criteria, and hold out these tasks as our uncertainty prediction tasks.
@@ H.1 Target Calibration Tasks
H.2 Instruction-Tuning Tasks
Because most modern instruction-tuned models are trained on all of Super-NaturalInstructions, they have seen the our calibration target tasks during instruction-tuning. Therefore, we instruction-tune our own T5 model to test the effectiveness of our method on unseen tasks. Our model is trained on a subset of the SuperNatural-Instructions dataset (wang-etal-2022-super). Specifically, we instruction-tune on the English split used in the original paper but we take out our target calibration tasks identified in § H.1. This gives us a training dataset of $879$ instruction-tuning tasks, with a total of roughly 1.2M training samples total.
@@ H.2 Instruction-Tuning Tasks
Appendix I Ministral plots
In Figure 6 we display the plots with Ministral-8B. As reflected in the AUROC score in Table 1, calibration is slightly worse; however, compared to baselines, it still does a more accurate job of forecasting accuracy.
@@ Appendix I Ministral plots
Appendix J Instruction-Tuning T5
@@
We follow a standard recipe for instruction-tuning T5-Large, established in wang-etal-2022-super. Specifically, we tune the model for 3 epochs with a batch size of $16$ and a learning rate of $5\times 10^{-3}$. We use the AdamW optimizer, and a constant learning rate schedule after a warmup period of $500$ steps. During instruction-tuning, we train the model with the semantic definition of each task prepended to the task input, and we similarly prompt the model when performing our target Q&A tasks.
Appendix K Resource Reporting
K.1 Compute Resources
@@
Here we report the compute resources used in this work. Instruction-tuning T5 took a total of $200$ GPU hours across $4$ NVIDIA-V100s. Running uncertainty distillation on Instruct-T5 and FLAN-T5 took 16 hours per model on a single NVIDIA-H100. Finetuning T5-base on SQUAD for our initial model took 3 hours on a single NVIDIA RTX 2080, and training using uncertainty distillation took 8 hours on a single NVIDIA-V100. Finetuning Ministral-8B (LoRA) and finetuning Llama-3B each took took three hours on two NVIDIA-A100s. Our lexical baseline for SQUAD took one hour on one NVIDIA RTX 2080; for SNI took three hours on one NVIDIA RTX 2080; for MMLU took three hours on one NVIDIA-A100; for SocialIQA took three hours on one NVIDIA-A100; for GSM8K took four hours on one NVIDIA-A100. Prompting for MMLU and prompting for SocialIQA took 1 hour on one NVIDIA-A100. Sampling for SQUAD took a total of 60 GPU hours on NVIDIA-V100s; for SNI took 45 GPU hours on NVIDIA-A100s; for SocialIQA took 350 hours on NVIDIA-A100s; for MMLU took 350 hours on NVIDIA-A100s; for GSM8k took 80 hours on NVIDIA-A100s.
K.2 Resource Intended Use
@@
Super-NaturalInstructions (SNI) is an open-source instruction tuning dataset, released under the Apache License. [^21] Available here: https://github.com/allenai/natural-instructions The intended use of SNI is to instruction-tune language models to learn to follow instructions, and to evaluate a model’s ability to follow instructions on unseen tasks. While we use the SNI dataset for precisely this purpose during instruction-tuning, we also use $15$ held-out tasks to serve as uncertainty quantification tasks. This does not necessarily fall under the intended use of instruction-tuning; however, the authors of SNI also mention that the dataset may serve as a large, multi-task natural language resource (wang-etal-2022-super), and our usage of the target calibration tasks does fall under this use case.
@@
The Stanford Question Answering Dataset (SQUAD) (rajpurkar2016squad) is distributed under the Creative Commons Attribution-Sharealike 4.0 license, which permits use of the dataset as long as it is properly attributed and as long as the results are distributed under the same license. As we cite the paper and plan to publically release our code and models after acceptance, our use of this dataset is permitted under this license.
@@
SocialIQA (siqa) is not explicitly licensed, but they state that they “ establish Social IQa as a resource” for future models.
@@
MMLU (mmlu), GSM8K(gsm8k), and MMLU-pro(wang2024mmluprorobustchallengingmultitask) are published under the MIT license, which allows users to freely copy, use, and change the licensed material.
Appendix L Uncertainty distillation hyperparameters
@@
Hyperparameters for fine-tuning via API
For the experiments reported in § 7, we fine-tune the gemini-2.5-flash-lite model using LoRA with rank $4$ for $10$ epochs and defaults for other hyperparameters. On validation data, we compared performance for different numbers of incorrect examples (§ 3.3), finding that augmenting the tuning set with a single incorrect prediction had marginal impact on accuracy while significantly improving calibration. We also compared both temperature scaling and isotonic regression, finding that isotonic scaling produced better calibration, while temperature scaling produced higher accuracy. To fit the calibration map, we held out 10% of the training data.
Model Epochs Learning rate Batch size Grad accum steps
@@
Table 11: Hyperparameters for training all T5 models but Instruct-T5 (see Appendix J for details). All models are trained with the AdamW optimizer.
@@
  $b\leftarrow$ bin(p)(p)
  $z\leftarrow$ verbalize_confidence_map(y^,b)($\hat${y},b)